# Graph-based Square-Root Estimation for Sparse Linear Regression

### v1

@@ 4.1 Simulation data
Table 6: Performance comparison of estimation, prediction and model selection for Example 3 with t distribution noise.
$L_{2}$ distance RPE FPR FNR
(I) 20/20/400 Ridge 9.225 (0.700) 7.231 (1.661) 0.960 (0.014) 0.007 (0.020)
Lasso 10.969 (1.269) 6.812 (2.440) 0.094 (0.018) 0.509 (0.112)
Alasso 10.993 (1.239) 6.464 (2.420) 0.086 (0.022) 0.476 (0.123)
Enet 8.842 (0.980) 6.792 (1.962) 0.416 (0.044) 0.125 (0.117)
SRL 10.737 (1.373) 6.420 (2.527) 0.173 (0.027) 0.383 (0.154)
SRIG-o 10.397 (0.390) 11.387 (1.191) 0.839 (0.098) 0.027 (0.075)
SRIG 9.823 (0.508) 9.980 (1.356) 0.915 (0.040) 0.016 (0.051)
GSRE-o 6.953 (1.613) 4.088 (2.599) 0.289 (0.102) 0.051 (0.107)
GSRE 6.177 (1.608) 3.240 (2.432) 0.317 (0.116) 0.023 (0.071)
(II) 40/40/400 Ridge 7.625 (0.606) 3.650 (0.952) 0.963 (0.013) 0.000 (0.000)
Lasso 7.635 (1.138) 2.832 (1.166) 0.050 (0.019) 0.193 (0.096)
Alasso 7.707 (1.349) 2.156 (1.001) 0.046 (0.019) 0.193 (0.103)
Enet 6.968 (0.793) 4.426 (1.238) 0.306 (0.049) 0.021 (0.044)
SRL 7.539 (1.224) 2.707 (1.168) 0.057 (0.024) 0.175 (0.096)
SRIG-o 10.262 (0.289) 11.230 (1.078) 0.496 (0.105) 0.008 (0.026)
SRIG 9.518 (0.550) 9.579 (1.412) 0.583 (0.146) 0.001 (0.009)
GSRE-o 5.626 (1.393) 2.881 (2.003) 0.039 (0.028) 0.015 (0.041)
GSRE 4.714 (1.007) 1.803 (1.089) 0.033 (0.026) 0.000 (0.000)
(III) 60/60/400 Ridge 6.925 (0.752) 2.341 (0.577) 0.962 (0.012) 0.000 (0.000)
Lasso 5.864 (1.131) 1.657 (0.742) 0.040 (0.023) 0.077 (0.081)
Alasso 6.138 (1.175) 1.205 (0.537) 0.026 (0.015) 0.087 (0.070)
Enet 6.873 (0.847) 4.703 (1.232) 0.188 (0.051) 0.013 (0.036)
SRL 5.737 (1.135) 1.600 (0.717) 0.044 (0.025) 0.075 (0.073)
SRIG-o 4.963 (0.866) 2.233 (0.905) 0.026 (0.022) 0.000 (0.000)
SRIG 4.188 (0.771) 1.439 (0.671) 0.040 (0.013) 0.000 (0.000)
GSRE-o 4.018 (0.813) 1.198 (0.590) 0.034 (0.025) 0.001 (0.009)
GSRE 3.782 (0.745) 1.060 (0.545) 0.030 (0.018) 0.000 (0.000)
@@
Table 7: Performance comparison of estimation, prediction and model selection for Example 3 with laplace distribution noise.
$L_{2}$ distance RPE FPR FNR
(I) 20/20/400 Ridge 9.188 (0.634) 7.307 (1.556) 0.962 (0.015) 0.007 (0.020)
Lasso 10.950 (1.347) 6.687 (2.558) 0.093 (0.022) 0.512 (0.118)
Alasso 10.911 (1.484) 6.328 (2.830) 0.086 (0.021) 0.465 (0.134)
Enet 8.800 (0.957) 6.712 (1.886) 0.401 (0.047) 0.103 (0.101)
SRL 10.647 (1.492) 6.193 (2.783) 0.170 (0.029) 0.399 (0.161)
SRIG-o 10.386 (0.337) 11.350 (1.039) 0.836 (0.068) 0.017 (0.038)
SRIG 9.809 (0.481) 9.939 (1.263) 0.909 (0.052) 0.011 (0.034)
GSRE-o 6.891 (1.680) 3.975 (2.583) 0.260 (0.104) 0.045 (0.101)
GSRE 6.048 (1.504) 3.015 (2.157) 0.299 (0.122) 0.016 (0.061)
(II) 40/40/400 Ridge 7.527 (0.634) 3.646 (1.012) 0.960 (0.017) 0.000 (0.000)
Lasso 7.398 (1.056) 2.685 (1.168) 0.049 (0.019) 0.185 (0.101)
Alasso 7.195 (1.161) 1.854 (0.800) 0.040 (0.019) 0.151 (0.100)
Enet 6.874 (0.815) 4.327 (1.202) 0.303 (0.056) 0.013 (0.033)
SRL 7.129 (0.957) 2.412 (1.005) 0.059 (0.026) 0.151 (0.091)
SRIG-o 10.259 (0.289) 11.221 (1.062) 0.504 (0.086) 0.011 (0.028)
SRIG 9.571 (0.467) 9.684 (1.265) 0.593 (0.108) 0.004 (0.021)
GSRE-o 5.359 (1.264) 2.543 (1.541) 0.035 (0.028) 0.012 (0.032)
GSRE 4.346 (1.185) 1.555 (1.197) 0.033 (0.028) 0.000 (0.000)
(III) 60/60/400 Ridge 6.645 (0.714) 2.170 (0.474) 0.959 (0.015) 0.000 (0.000)
Lasso 5.479 (1.159) 1.478 (0.677) 0.041 (0.019) 0.063 (0.078)
Alasso 5.654 (1.263) 1.037 (0.450) 0.024 (0.014) 0.071 (0.065)
Enet 6.669 (0.931) 4.419 (1.363) 0.187 (0.047) 0.005 (0.018)
SRL 5.326 (1.072) 1.353 (0.560) 0.047 (0.026) 0.047 (0.062)
SRIG-o 4.866 (0.716) 2.115 (0.753) 0.025 (0.016) 0.000 (0.000)
SRIG 3.999 (0.605) 1.280 (0.538) 0.041 (0.008) 0.000 (0.000)
GSRE-o 3.750 (0.671) 1.016 (0.450) 0.034 (0.023) 0.000 (0.000)
GSRE 3.524 (0.639) 0.890 (0.428) 0.028 (0.017) 0.000 (0.000)

### v2

@@ 4.1 Simulation data
Table 4: Performance comparison of estimation, prediction and model selection for Example 3 with uniform noise.
$L_{2}$ distance RPE FPR FNR MCC
(I) 20/20/400 Ridge 8.127(1.315) 4.979(2.447) 0.952(0.019) 0.007(0.024) 0.074(0.035)
Lasso 10.704(1.523) 6.192(2.880) 0.075(0.021) 0.487(0.171) 0.445(0.173)
Alasso 11.246(1.370) 6.429(2.626) 0.054(0.020) 0.527(0.141) 0.465(0.147)
Enet 8.245(1.245) 3.518(2.011) 0.143(0.029) 0.185(0.117) 0.560(0.113)
SRL 11.241(1.350) 6.724(2.731) 0.117(0.030) 0.484(0.124) 0.375(0.130)
SRIG-o 10.393(0.337) 11.387(1.015) 0.842(0.065) 0.024(0.052) 0.138(0.059)
SRIG 10.393(0.337) 11.387(1.015) 0.842(0.065) 0.024(0.052) 0.138(0.059)
GSRE-o 7.045(1.609) 4.483(2.545) 0.312(0.119) 0.039(0.080) 0.484(0.120)
GSRE 7.045(1.609) 4.483(2.545) 0.312(0.119) 0.039(0.080) 0.484(0.120)
(II) 40/40/400 Ridge 7.721(0.639) 3.827(1.099) 0.955(0.013) 0.000(0.000) 0.083(0.013)
Lasso 8.156(1.058) 3.784(1.538) 0.029(0.018) 0.285(0.121) 0.725(0.108)
Alasso 7.710(1.191) 2.017(0.872) 0.034(0.020) 0.204(0.101) 0.767(0.102)
Enet 6.757(0.973) 3.916(1.457) 0.117(0.025) 0.064(0.070) 0.686(0.075)
SRL 8.210(1.197) 4.045(1.749) 0.026(0.019) 0.303(0.147) 0.722(0.131)
SRIG-o 7.713(1.534) 6.256(2.657) 0.342(0.203) 0.000(0.000) 0.525(0.219)
SRIG 7.713(1.534) 6.256(2.657) 0.342(0.203) 0.000(0.000) 0.525(0.219)
GSRE-o 5.059(1.195) 2.119(1.400) 0.039(0.035) 0.011(0.034) 0.890(0.077)
GSRE 5.059(1.195) 2.119(1.400) 0.039(0.035) 0.011(0.034) 0.890(0.077)
(III) 60/60/400 Ridge 6.685(0.683) 2.366(0.633) 0.958(0.014) 0.000(0.000) 0.080(0.013)
Lasso 6.577(1.508) 2.670(1.777) 0.017(0.016) 0.165(0.156) 0.845(0.090)
Alasso 6.645(1.177) 1.256(0.507) 0.024(0.016) 0.139(0.084) 0.842(0.068)
Enet 6.353(1.085) 3.776(1.478) 0.053(0.026) 0.045(0.064) 0.829(0.057)
SRL 6.123(1.181) 2.054(0.972) 0.023(0.015) 0.103(0.090) 0.866(0.058)
SRIG-o 4.874(0.794) 2.123(0.844) 0.021(0.010) 0.001(0.009) 0.936(0.026)
SRIG 4.874(0.794) 2.123(0.844) 0.021(0.010) 0.001(0.009) 0.936(0.026)
GSRE-o 4.170(0.907) 1.381(0.769) 0.017(0.017) 0.001(0.009) 0.948(0.045)
GSRE 4.170(0.907) 1.381(0.769) 0.017(0.017) 0.001(0.009) 0.948(0.045)
Besides these three examples, we also provide a complex simulation example and conduct a sensitivity analysis of Assumption 2.2 in the supplementary materials. The numerical results indicate that our proposed GSRE method still outperforms the other methods. In summary, the simulation results indicate that the proposed GSRE method, which combines square-root estimation with graph penalty, shows superior performance in estimation, prediction and model selection under various noise environments. This result highlights the general applicability and reliability of the GSRE method in statistical modeling.
{sidewaystable}
@@
[h] Performance comparison of estimation, prediction and model selection for Example 1&2 with Gaussian noise. Example 1 Example 2 $L_{2}$ distance RPE FPR FNR MCC $L_{2}$ distance RPE FPR FNR MCC (I) 20/20/400 Ridge 6.833(1.252) 3.050(3.099) 0.948(0.020) 0.001(0.009) 0.087(0.020) 6.820(1.356) 3.424(2.814) 0.951(0.016) 0.000(0.000) 0.086(0.015) Lasso 12.841(1.684) 3.426(2.338) 0.061(0.024) 0.508(0.105) 0.465(0.128) 11.335(1.217) 5.704(3.023) 0.071(0.023) 0.497(0.117) 0.451(0.131) Alasso 13.044(2.072) 3.334(2.795) 0.048(0.024) 0.497(0.117) 0.510(0.144) 11.449(1.490) 5.155(2.982) 0.048(0.021) 0.476(0.134) 0.523(0.143) Enet 9.317(1.443) 2.473(2.346) 0.150(0.027) 0.265(0.105) 0.491(0.101) 8.581(1.239) 3.164(2.006) 0.143(0.033) 0.217(0.095) 0.538(0.110) SRL 12.550(1.702) 3.838(2.701) 0.123(0.023) 0.447(0.109) 0.394(0.106) 11.571(1.403) 5.708(2.832) 0.130(0.021) 0.440(0.114) 0.388(0.117) SRIG-o 8.819(0.661) 13.752(2.380) 0.576(0.061) 0.000(0.000) 0.317(0.036) 9.235(0.516) 12.075(1.678) 0.562(0.058) 0.011(0.049) 0.316(0.046) SRIG 8.819(0.661) 13.752(2.380) 0.576(0.061) 0.000(0.000) 0.317(0.036) 9.233(0.519) 12.071(1.682) 0.562(0.058) 0.011(0.049) 0.316(0.046) GSRE-o 3.219(1.057) 1.392(1.772) 0.128(0.048) 0.000(0.000) 0.718(0.074) 4.319(1.165) 1.941(2.111) 0.160(0.051) 0.007(0.047) 0.666(0.070) GSRE 3.219(1.057) 1.392(1.772) 0.128(0.048) 0.000(0.000) 0.718(0.074) 4.320(1.164) 1.941(2.111) 0.160(0.051) 0.007(0.047) 0.666(0.071) (II) 40/40/400 Ridge 6.325(0.695) 2.078(0.929) 0.966(0.018) 0.000(0.000) 0.071(0.019) 6.813(0.879) 2.864(1.108) 0.962(0.015) 0.000(0.000) 0.076(0.015) Lasso 9.765(1.470) 2.401(1.123) 0.019(0.015) 0.353(0.110) 0.709(0.088) 8.063(1.345) 2.946(1.329) 0.022(0.017) 0.255(0.130) 0.768(0.105) Alasso 9.758(1.469) 0.984(0.440) 0.024(0.017) 0.268(0.099) 0.751(0.096) 8.184(1.392) 1.603(0.750) 0.025(0.016) 0.213(0.105) 0.787(0.093) Enet 4.696(0.852) 2.998(1.419) 0.105(0.031) 0.000(0.000) 0.753(0.053) 5.668(0.913) 3.400(1.454) 0.118(0.030) 0.019(0.049) 0.718(0.066) SRL 9.554(1.531) 2.459(1.392) 0.016(0.014) 0.333(0.116) 0.732(0.097) 7.970(1.372) 3.042(1.553) 0.024(0.015) 0.249(0.114) 0.766(0.090) SRIG-o 3.300(1.290) 2.000(1.967) 0.038(0.067) 0.000(0.000) 0.914(0.124) 3.944(1.283) 2.144(1.927) 0.045(0.067) 0.000(0.000) 0.896(0.127) SRIG 3.300(1.290) 2.000(1.967) 0.038(0.067) 0.000(0.000) 0.914(0.124) 3.945(1.282) 2.142(1.928) 0.045(0.067) 0.000(0.000) 0.896(0.127) GSRE-o 2.431(0.616) 0.751(0.598) 0.009(0.010) 0.000(0.000) 0.972(0.029) 3.011(0.613) 0.811(0.524) 0.014(0.011) 0.000(0.000) 0.956(0.031) GSRE 2.431(0.616) 0.751(0.598) 0.009(0.010) 0.000(0.000) 0.972(0.029) 3.014(0.613) 0.811(0.524) 0.014(0.011) 0.000(0.000) 0.956(0.031) (III) 60/60/400 Ridge 6.267(0.781) 2.272(0.605) 0.957(0.014) 0.000(0.000) 0.081(0.014) 6.348(0.777) 2.285(0.596) 0.956(0.013) 0.000(0.000) 0.082(0.012) Lasso 8.110(1.492) 1.497(0.941) 0.010(0.010) 0.223(0.121) 0.829(0.081) 6.211(1.549) 1.725(1.214) 0.009(0.011) 0.108(0.135) 0.906(0.082) Alasso 9.619(1.312) 0.736(0.230) 0.020(0.012) 0.261(0.072) 0.772(0.061) 7.092(1.272) 1.042(0.418) 0.018(0.014) 0.135(0.079) 0.862(0.069) Enet 3.903(0.613) 2.137(0.856) 0.055(0.030) 0.000(0.000) 0.855(0.069) 4.715(1.023) 2.579(1.403) 0.047(0.029) 0.011(0.037) 0.865(0.076) SRL 7.791(1.332) 1.248(0.615) 0.013(0.012) 0.172(0.105) 0.851(0.075) 5.931(1.453) 1.489(0.998) 0.013(0.013) 0.072(0.099) 0.916(0.074) SRIG-o 2.269(0.446) 0.734(0.385) 0.004(0.007) 0.000(0.000) 0.987(0.022) 2.701(0.583) 0.792(0.461) 0.001(0.004) 0.000(0.000) 0.996(0.011) SRIG 2.269(0.446) 0.734(0.385) 0.004(0.007) 0.000(0.000) 0.987(0.022) 2.706(0.582) 0.792(0.461) 0.001(0.004) 0.000(0.000) 0.996(0.011) GSRE-o 2.028(0.341) 0.413(0.241) 0.005(0.007) 0.000(0.000) 0.984(0.022) 2.464(0.477) 0.511(0.306) 0.003(0.006) 0.000(0.000) 0.991(0.019) GSRE 2.028(0.341) 0.413(0.241) 0.005(0.007) 0.000(0.000) 0.984(0.022) 2.468(0.476) 0.511(0.306) 0.003(0.006) 0.000(0.000) 0.991(0.019)
{sidewaystable}